The function f: R^3 to R , defined by

f(x,y,z)= x^3+ e^(y+z) is differentiable everywhere on R^3.

True or false with full explanation


1
Expert's answer
2021-06-22T10:41:58-0400

Taking into account that the elementary functions "g(x,y,z)=x^3", "h(x,y,z)=y+z" and "l(x,y,z)=e^x" are differentiable everywhere on "\R^3", we conclude that their composition "f(x,y,z)= x^3+ e^{y+z}" is differentiable everywhere on "\R^3".


Answer: true